OpenStudy (anonymous):

Calculate for following integral : \[\int\limits cosecx dx\] by substituting t=tan(x/2)

OpenStudy (mimi_x3):

\[ \int \frac{1}{sin(x)}dx\] \[ \int \frac{1}{ \frac{2t}{1+t^2}}dt = \int \frac{1+t^2}{2t}dt = \int \frac{1}{2t} + \frac{t^2}{2t} dt\]

OpenStudy (mimi_x3):

\[ \frac{1}{2} \int \frac{1}{t} dt + \int \frac{t}{2}dt\]

OpenStudy (mimi_x3):

\[\frac{1}{2} \int \frac{1}{t} dt + \frac{1}{2}\int tdt\]

OpenStudy (anonymous):

Ok, so then you get 1/2lnt + 1/4t^2?

OpenStudy (mimi_x3):

hmm..something is wrong.

OpenStudy (anonymous):

with my answer?

OpenStudy (mimi_x3):

with mine..

OpenStudy (anonymous):

I think I know, we should have substituted 2dt/1+t\[^2\] in for dx, as it says on the wiki page

OpenStudy (mimi_x3):

im sooo stupiddd lol! i forgot about \[ dx = \frac{2}{1+t}\] lol

OpenStudy (anonymous):

haha don't worry, I've followed through with that and it works out

OpenStudy (mimi_x3):

\[ \int \frac{1}{sinx} dx = \int \frac{ 1+t^2}{2t} * \frac{2}{1+t^2}\] \[ \int \frac{1}{t} dt = ln|t| +c\] better now hehe:D

OpenStudy (anonymous):

yep exactly